OpenStudy (just_one_last_goodbye):

The legs of a right triangle are 5m and 12m long. What is the length of the hypotenuse? Explain how you found your answer.

OpenStudy (anonymous):

13

OpenStudy (just_one_last_goodbye):

ok and what were the steps?

OpenStudy (anonymous):

x^2 = (5)^2 + (12)^2 by c^2 = a^2 + b^2

OpenStudy (math&ing001):

Use Pythagoras formula : H²=A²+B²

OpenStudy (anonymous):

x^2 = 25+144 x^2 = 169 x = 13

OpenStudy (just_one_last_goodbye):

hmm you seem like u know what your doing my friend ^_^

OpenStudy (anonymous):

exactly :)

OpenStudy (anonymous):

just trying to sharp ma mathematical skills!

OpenStudy (just_one_last_goodbye):

lol like sharpening a pencil ;)just dont go stabbing people with the point